# Self-Supervised Multi-Part Articulated Objects Modeling via Deformable Gaussian Splatting and Progressive Primitive Segmentation

### v1

@@ D.2 Result on the RS-Art Dataset
Object Method CD-$\downarrow$ CD-m 0$\downarrow$ CD-m 1$\downarrow$ CD-$\downarrow$
@@
D.3 Ablation Studies
Quantitative impact of the refinement stage.
Sec. 5.5 in the main paper visualizes the qualitative benefit of coarse to fine refinement. Tab. 5 complements that visual evidence with numerical results on four representative objects. Across all geometry metrics, the refined model (w/ Refined) consistently outperforms the same network trained without the refinement stage (w/o Refined). Averaged across the four objects, the refinement reduces CD-s by $35\%$ and CD-m by $60\%$. The greatest gains occur at part boundaries (CD-m columns), confirming that boundary-aware Gaussian splitting and local fine-tuning markedly sharpen the reconstructed surfaces and prevent inter-penetration. These numbers corroborate the qualitative observations in Sec. 5.5 and highlight the refinement stage as a key component for accurate part-level geometry.
@@ D.3 Ablation Studies
Impact of the number of interaction states.
Tab. 6 and Fig. 12 show how our method reacts to different numbers of input interaction states $K$. With just two interaction states ($K=2$), the network sees only a start-to-end displacement for each Gaussian. This limited evidence allows multiple trajectory assignments to fit the same images, so Gaussians can drift between neighboring parts, leading to inflated joint errors and indistinct part boundaries. As more states are provided, every primitive must follow a longer and more distinctive trajectory. The stronger constraint eliminates ambiguity: a Gaussian that does not truly belong to a part can no longer match the motion of that part in all states, and the optimizer moves it to the correct rigid component, causing a rapid drop in error. However, after each joint has appeared in several diverse poses, additional states add little fresh kinematic information while still introducing measurement noise and extra training time, so the accuracy curve levels off and may fluctuate slightly.
@@ Appendix E Limitations
Incomplete part reconstruction under occlusion.
Like other NeRF- and 3D Gaussian Splatting-based approaches, DeGSS may struggle to reconstruct complete geometry and texture for every part of the object. This limitation arises in cases of severe occlusion, limited viewpoints, or coupled articulation, where certain surfaces are never observed. In addition, the representation is not surface-aware by design, with both the NeRF and 3DGS models radiance and density fields rather than explicit meshes—which can lead to surface artifacts and coarse or noisy geometry. Real-scene results also suggest that texture fidelity can degrade under challenging lighting conditions. Future work may explore integrating surface priors or reflectance models to improve reconstruction quality in under-constrained regions.
Limited articulation understanding from sparse states.
DeGSS infers motion by comparing interaction states, but reasoning about joint types and motion limits remains difficult. In particular, screw-type joints are difficult to identify, as the motion between observed states may produce little or no visible appearance change, making them indistinguishable from static or simpler joints under visual comparison alone. Similarly, estimating articulation limits (e.g., joint range or end stops) requires observing motions near those extremes, which sparse or biased sampling may not capture. Incorporating additional interaction sequences or learning from temporal transitions could help enrich articulation inference in future extensions.
Lack of physical properties for simulation.
While DeGSS focuses on visual and kinematic modeling, it does not capture physical properties such as mass, friction, or compliance, which are essential for high-fidelity digital twins. These properties are crucial for accurate simulation and downstream robotic tasks, but are currently absent from most NeRF/3DGS-based pipelines. Bridging this gap may require integrating material prediction models, learning from physical interactions, or coupling neural reconstruction with differentiable simulation frameworks.

### v2

@@ D.7 Ablation Studies
Object Instance Method CD-s $\downarrow$ CD-m 0 $\downarrow$ CD-m 1 $\downarrow$ CD-w $\downarrow$
@@
Refinement stage. Sec. 5.5 in the main paper visualizes the qualitative benefit of coarse-to-fine refinement. Tab. 11 complements that visual evidence with numerical results on four representative objects. Across all geometry metrics, the refined model (w/ Refined) consistently outperforms the same network trained without the refinement stage (w/o Refined). Averaged across the four objects, the refinement reduces CD-s by $35\%$ and CD-m by $60\%$. The greatest gains occur at part boundaries (CD-m columns), confirming that boundary-aware Gaussian splitting and local fine-tuning markedly sharpen the reconstructed surfaces and prevent inter-penetration. These numbers corroborate the qualitative observations in Sec. 5.5 and highlight the refinement stage as a key component for accurate part-level geometry.
@@
The number of interaction states. Tab. 12 and Fig. 16 show how our method reacts to different numbers of input interaction states $K$. With just two interaction states ($K=2$), the network sees only a start-to-end displacement for each Gaussian. This limited evidence allows multiple trajectory assignments to fit the same images, so Gaussians can drift between neighboring parts, leading to inflated joint errors and indistinct part boundaries. As more states are provided, every primitive must follow a longer and more distinctive trajectory. The stronger constraint eliminates ambiguity: a Gaussian that does not truly belong to a part can no longer match the motion of that part in all states, and the optimizer moves it to the correct rigid component, causing a rapid drop in error. However, after each joint has appeared in several diverse poses, additional states add little fresh kinematic information while still introducing measurement noise and extra training time, so the accuracy curve levels off and may fluctuate slightly.
Table 13: Movable-component count estimation accuracy (correct/total number) for three VLM prompting schemes across eight object categories (instance counts in parentheses).
Method Box (2) Door (3) Eyeglasses (2) Faucet (2) Oven (2) Fridge (3) Storage (3) Table (3) Overall Accuracy
VLM‑M1 1 / 2 3 / 3 2 / 2 1 / 2 2 / 2 3 / 3 3 / 3 3 / 3 90 %
VLM‑M2 0 / 2 0 / 3 1 / 2 2 / 2 0 / 2 0 / 3 0 / 3 0 / 3 15 %
VLM‑Ours 2 / 2 3 / 3 2 / 2 2 / 2 2 / 2 3 / 3 3 / 3 3 / 3 100 %
Prompting strategies on component-count estimation. Tab. 13 compares our approach with two alternative prompting strategies, specifically single-view queries and generic cross-state queries, and shows that our method provides the most stable and accurate component-count estimates. VLM-M1 queries single views (“How many movable components does this object contain?”); VLM-M2 queries cross-state pairs (“How many movable parts are identifiable from these two states of the same object?”). Final counts are determined by the most frequent response. This design eliminates the need for manual $n_{parts}$ selection required in previous 3DGS pipelines.
Table 14: Comparison of SAM-based segmentation accuracy between our approach and the SAGD method.
Metric Method Box Door Eyeglasses Faucet Oven Fridge Storage Table Mean
IoU SAGD (Hu et al., 2024) 45.37 42.79 28.49 34.71 38.32 45.02 44.70 58.04 42.18
SAM-Ours 84.20 89.61 69.40 83.72 75.33 87.09 78.06 79.25 80.83
Acc SAGD (Hu et al., 2024) 94.63 97.14 97.96 97.46 91.64 95.23 94.54 97.79 95.80
SAM-Ours 98.06 98.75 99.57 99.47 98.70 99.37 97.92 98.76 98.82
F1-score SAGD (Hu et al., 2024) 54.46 52.34 38.01 49.02 46.89 52.11 55.31 65.21 51.67
SAM-Ours 89.74 95.05 81.04 90.82 81.42 91.46 85.00 85.58 87.51
SAM-based component segmentation. For SAM segmentation, as shown in Tab. 14, we compare our approach to the current state-of-the-art point-prompt SAM-based 3DGS segmentation method, SAGD (Hu et al., 2024). In all cases, these alternatives perform worse than the full pipeline, providing quantitative evidence for the value of each component.
Table 15: Quantitative evaluation of each pipeline module on all object categories. All metrics are averaged over parts. For Acc and IoU (reported as %), higher values are better; for CD-w (reported in $m$), lower values are better.
Module & Metric Box Door Eyeglasses Faucet Oven Fridge Storage Table Mean
Deformable GS (CD-w) 8.32 0.64 0.27 0.57 13.78 9.43 3.94 3.05 5.00
Coarse Seg. (Acc, %) 72.63 86.93 86.45 66.82 59.45 67.57 67.64 82.45 73.74
SAM+Prompt (IoU, %) 84.20 89.61 69.40 83.72 75.33 87.09 78.06 79.25 80.83
Refined Seg. (Acc,%) 95.83 95.41 98.64 99.53 99.23 96.41 85.76 98.29 96.14
Each pipeline module. For each key module, we report a dedicated quantitative metric in the Tab. 15. For Deformable GS reconstruction(Sec. 3.2), Chamfer distance demonstrates stable and accurate geometric modeling across all categories. For VLM-based part counting (Sec. 3.3), Delivers consistently high accuracy, validating the effectiveness of our task-specific prompt. For Coarse Gaussian Segmentation (Sec. 3.3), High accuracy confirms reliable part-level grouping, enabling robust downstream point prompt for SAM. For SAM with prompt (Sec. 3.4), Achieves precise semantic segmentation. For Refined segmentation(Sec. 3.4), Post-refinement clustering further increases accuracy, showing precise decoupling of part-level Gaussians. The consistency of these metrics across all objects and scenes confirms that each module is robust and that the pipeline as a whole is stable.
@@ Appendix E Limitations
Incomplete part reconstruction under occlusion. Like other NeRF-based and 3D Gaussian Splatting-based approaches, PD2GS may struggle to reconstruct complete geometry and texture for every part of the object. This limitation arises in cases of severe occlusion, limited viewpoints, or coupled articulation, where certain surfaces are never observed. In addition, the representation is not surface-aware by design, with both the NeRF and 3DGS models radiance and density fields rather than explicit meshes, which can lead to surface artifacts and coarse or noisy geometry. Real-scene results also suggest that texture fidelity can degrade under challenging lighting conditions. Future work may explore integrating surface priors, reflectance models, or learned data-driven priors, such as diffusion models, to enhance reconstruction fidelity and plausibly complete unobserved geometry in under-constrained regions.
Limited articulation understanding from sparse states. PD2GS infers motion by comparing interaction states, but reasoning about joint types and motion limits remains difficult. In particular, screw-type joints are difficult to identify, as the motion between observed states may produce little or no visible appearance change, making them indistinguishable from static or simpler joints under visual comparison alone. Similarly, estimating articulation limits (e.g., joint range or end stops) requires observing motions near those extremes, which sparse or biased sampling may not capture. Incorporating additional interaction sequences or learning from temporal transitions could help enrich articulation inference in future extensions.
Lack of physical properties for simulation. While PD2GS focuses on visual and kinematic modeling, it does not capture physical properties such as mass, friction, or compliance, which are essential for high-fidelity digital twins. These properties are crucial for accurate simulation and downstream robotic tasks, but are currently absent from most NeRF/3DGS-based pipelines. Bridging this gap may require integrating material prediction models, learning from physical interactions, or coupling neural reconstruction with differentiable simulation frameworks.